# How Far Are We on the Decision-Making of LLMs? Evaluating LLMs’ Gaming Ability in Multi-Agent Environments

### v3

@@ Abstract
Decision-making, a complicated task requiring various types of abilities, presents an excellent framework for assessing Large Language Models (LLMs). Our research investigates decision-making capabilities of LLMs through the lens of Game Theory. We focus specifically on games that support the simultaneous participation of more than two agents. We introduce GAMA($\gamma$)-Bench, which evaluates LLMs’ Gaming Ability in Multi-Agent environments. $\gamma$-Bench includes eight classical multi-agent games and a scoring scheme specially designed to quantitatively assess LLMs’ performance. Leveraging $\gamma$-Bench, we investigate LLMs’ robustness, generalizability, and strategies for enhancement. Results reveal that while GPT-3.5 shows satisfying robustness, its generalizability is relatively limited. However, its performance can be improved through approaches such as Chain-of-Thought. Additionally, we evaluate twelve versions from six models, including GPT-3.5, GPT-4, Gemini, LLaMA-3.1, Mixtral, and Qwen-2. We find that Gemini-1.5-Pro outperforms other models with a score of $63.8$ out of $100$, followed by LLaMA-3.1-70B and GPT-4 with scores of $60.9$ and $60.5$, respectively. The code and experimental results are made publicly available via https://github.com/CUHK-ARISE/GAMABench.
Figure 1: $\gamma$-Bench enables multiple LLMs and humans to engage in multi-round games. The framework comprises three categories of games, each targeting different LLM abilities, and includes eight classic games from Game Theory.
@@ 1 Introduction
We have recently witnessed the advancements in Artificial Intelligence (AI) made by Large Language Models (LLMs), which have marked a significant breakthrough in the field. ChatGPT [^1] https://chat.openai.com/, a leading LLM, has demonstrated its proficiency in a variety of Natural Language Processing (NLP) tasks, including machine translation Jiao et al. (2023), sentence revision Wu et al. (2023a), information retrieval Zhu et al. (2023), and program repair Surameery & Shakor (2023). Beyond the academic sphere, LLMs have entered diverse aspects of our everyday life, such as education Baidoo-Anu & Ansah (2023), legal service Guha et al. (2023), product design Lanzi & Loiacono (2023), and healthcare Johnson et al. (2023). Given their extensive capabilities, evaluating LLMs demands more than simple, isolated tasks. A comprehensive and multifaceted approach is highly in demand to assess the efficacy of these advanced models.
With the broad knowledge encoded in LLMs, their intelligence Liang et al. (2023b), and capabilities in general-purpose task solving Qin et al. (2023), a question emerges: can LLMs assist in everyday decision-making? Decision-making is a complex task requiring various abilities: (1) Perception: the ability to understand situations, environments, and rules, and extends to long-text understanding for LLMs. (2) Planning: the ability to maximize long-term over immediate benefits by anticipating potential outcomes. (3) Arithmetic Reasoning: the ability to quantify real-world scenarios and perform calculations. (4) ToM Reasoning: the Theory of Mind Kosinski (2023); Bubeck et al. (2023) refers to the ability to infer others’ intentions and beliefs. (5) Critical Thinking: the ability to integrate all available information to arrive at the best decision. Given these complexities, decision-making poses a significant challenge for intelligent agents.
To develop a framework to evaluate the decision-making ability of LLMs, we draw upon the principles of Game Theory, a well-established field with extensive applicability. Our rationale is threefold: (1) Scope: Game theory allows for the abstraction of diverse real-life scenarios into simple mathematical models, facilitating a broad range of evaluations. (2) Quantifiability: By examining the Nash equilibrium Nash (1950) within these models, we gain a measurable metric for comparing LLMs’ decision-making performance. (3) Variability: The adjustable parameters of these models enable the creation of variant scenarios, enhancing the diversity and robustness of our assessments. Our framework evaluates LLMs in complex settings involving multi-player, multi-action, and multi-round games, incorporating eight classical games extensively analyzed in game theory literature.
Table 1: Performance (scores) of different closed-source LLMs on $\gamma$-Bench.
@@
0613 1106 0125 0125 1.0 1.5
Guess 2/3 of the Average $41.4_{\pm 0.5}$ $68.5_{\pm 0.5}$ $63.4_{\pm 3.4}$ $91.6_{\pm 0.6}$ $77.3_{\pm 6.2}$ $95.4_{\pm 0.5}$
El Farol Bar $74.8_{\pm 4.5}$ $64.3_{\pm 3.1}$ $68.7_{\pm 2.7}$ $23.0_{\pm 8.1}$ $33.5_{\pm 10.3}$ $37.2_{\pm 4.2}$
Divide the Dollar $42.4_{\pm 7.7}$ $70.3_{\pm 3.3}$ $68.6_{\pm 2.4}$ $98.1_{\pm 1.9}$ $77.6_{\pm 3.6}$ $93.8_{\pm 0.3}$
Public Goods Game $38.8_{\pm 8.1}$ $43.5_{\pm 12.6}$ $17.8_{\pm 1.7}$ $89.2_{\pm 1.8}$ $68.5_{\pm 7.6}$ $100.0_{\pm 0.0}$
Diner’s Dilemma $67.0_{\pm 4.9}$ $1.4_{\pm 1.3}$ $2.8_{\pm 2.8}$ $0.9_{\pm 0.7}$ $3.1_{\pm 1.5}$ $35.9_{\pm 5.3}$
Sealed-Bid Auction $5.5_{\pm 1.0}$ $3.4_{\pm 1.0}$ $4.2_{\pm 0.3}$ $8.7_{\pm 0.7}$ $14.9_{\pm 6.6}$ $13.3_{\pm 4.6}$
Battle Royale $19.5_{\pm 7.7}$ $35.7_{\pm 6.9}$ $28.6_{\pm 11.0}$ $86.8_{\pm 9.7}$ $16.5_{\pm 6.9}$ $81.3_{\pm 7.7}$
Pirate Game $68.4_{\pm 20.0}$ $69.6_{\pm 14.7}$ $71.6_{\pm 7.6}$ $85.4_{\pm 8.6}$ $57.4_{\pm 14.3}$ $53.7_{\pm 1.2}$
Overall $43.5_{\pm 2.2}$ $44.8_{\pm 3.6}$ $41.9_{\pm 2.3}$ $60.5_{\pm 2.7}$ $43.3_{\pm 3.2}$ $63.8_{\pm 1.6}$
Table 2: Performance (scores) of different open-source LLMs on $\gamma$-Bench.
@@
Public Goods Game $19.6_{\pm 1.0}$ $90.6_{\pm 3.6}$ $96.9_{\pm 0.8}$ $27.7_{\pm 11.7}$ $83.7_{\pm 3.5}$ $81.3_{\pm 5.9}$
@@
Sealed-Bid Auction $16.9_{\pm 1.8}$ $4.5_{\pm 0.7}$ $4.2_{\pm 1.2}$ $0.8_{\pm 0.4}$ $5.2_{\pm 1.8}$ $0.9_{\pm 0.2}$
Battle Royale $35.9_{\pm 12.1}$ $77.7_{\pm 26.0}$ $92.7_{\pm 10.1}$ $12.6_{\pm 9.5}$ $36.0_{\pm 21.0}$ $81.7_{\pm 9.6}$
Pirate Game $42.3_{\pm 7.8}$ $35.7_{\pm 7.3}$ $38.2._{\pm 10.8}$ $31.8_{\pm 5.0}$ $53.0_{\pm 2.6}$ $53.0_{\pm 1.1}$
Overall $48.9_{\pm 2.9}$ $60.9_{\pm 2.4}$ $57.0_{\pm 4.2}$ $38.6_{\pm 2.3}$ $57.5_{\pm 2.1}$ $52.4_{\pm 3.4}$
@@
We include eights games and divide them into three categories based on their characteristics. The first category in our framework evaluates LLMs’ ability to make optimal decisions by understanding game rules and recognizing patterns in other players’ behavior. A distinctive characteristic of these games is that individual players cannot achieve higher gains without cooperation, provided that other participants cooperate. Essentially, these games’ Nash equilibrium aligns with maximizing overall social welfare. We name such games as Cooperative Games, including (1) Guess 2/3 of the Average, (2) El Farol Bar, and (3) Divide the Dollar. The second category assesses the propensity of LLMs to prioritize self-interest, potentially betraying others for greater gains. In contrast to the first category, games in this category incentivize higher rewards for participants who betray their cooperative counterparts. Typically, the Nash equilibrium in these games leads to reduced social welfare. This category is termed Betraying Games, including (4) Public Goods Game, (5) Diner’s Dilemma, (6) Sealed-Bid Auction. Last but not least, we focus specifically on two games characterized by sequential decision-making processes, distinguishing them from the previous six games based on simultaneous decision-making. Sequential Games are the (7) Battle Royale and (8) Pirate Game.
@@
In this paper, we instruct ten agents, based on the GPT-3.5 (0125) model, to engage in the eight games, followed by an analysis of the results obtained. Subsequently, we assess the model’s robustness against multiple runs, temperature parameter alterations, and prompt template variations. Further exploration is conducted to ascertain if instructional prompts, such as Chain-of-Thought (CoT) Kojima et al. (2022), enhance the model’s decision-making capabilities. Additionally, the model’s capacity to generalize across diverse game settings is examined. Finally, we evaluate the performance of twelve LLMs, including GPT-3.5-Turbo (0613, 1106, 0125) OpenAI (2022), GPT-4-Turbo (0125) OpenAI (2023), Gemini-1.0-Pro Pichai & Hassabis (2023), Gemini-1.5-Pro Pichai & Hassabis (2024), LLaMA-3.1 (8B, 70B, 405B) Dubey et al. (2024), Mixtral (8x7B, 8x22B) Jiang et al. (2024), and Qwen-2-72B Yang et al. (2024).
The contribution of this paper can be summarized as:
  • This paper reviews and compares existing literature on the evaluation of LLMs using game theory models, emphasizing focus on multi-player setting and LLMs’ generalizability.
  • We introduce a novel evaluation perspective of LLMs, i.e., their Gaming Ability in Multi-Agent environments, and present our designed framework, GAMA($\gamma$)-Bench.
@@
  • The $\gamma$-Bench framework is applied to twelve LLMs, facilitating an in-depth analysis of their performance in multi-agent gaming scenarios.
2 Games in GAMA-Bench
@@
We collect eight games well studied in Game Theory and propose $\gamma$-Bench, a framework with multi-player, multi-round, and multi-action settings. Notably, $\gamma$-Bench allows the simultaneous participation of both LLMs and humans, enabling us to evaluate LLMs’ performance when playing against humans or fixed strategies. The remaining part of this section details each game.
@@ 2.1 Cooperative Games
Initially introduced by Ledoux (1981), the game involves players independently selecting an integer between 0 and 100 (inclusive). The winner is the player(s) choosing the number closest to two-thirds of the group’s average. A typical initial strategy might lead players to assume an average of 50, suggesting a winning number around $50\times\frac{2}{3}\approx 33$. However, if all participants adopt this reasoning, the average shifts to 33, thereby altering the winning number to approximately 22. The game has a PSNE where all players selecting zero results in a collective win.
@@
Proposed by Arthur (1994) and Huberman (1988), this game requires players to decide to either visit a bar for entertainment or stay home without communication. The bar, however, has a limited capacity and can only accommodate part of the population. In a classical scenario, the bar becomes overcrowded and less enjoyable if more than 60% of the population decides to go there. Conversely, if 60% or fewer people are present, the experience is more enjoyable than staying home. Imagine that if everyone adopts the same pure strategy, i.e., either everyone going to the bar or everyone staying home, then the social welfare is not maximized. Notably, the game lacks a PSNE but presents an MSNE, where the optimal strategy involves going to the bar with a 60% probability and staying home with a 40% probability.
@@ 2.3 Sequential Games
3 Vanilla Experiments
@@
Key Findings: The model’s decisions are mainly influenced by the outcomes of the preceding round rather than deriving from the reasoning of the optimal strategy. Although initially demonstrating suboptimal performance, the model can learn from historical data and enhance its performance over time. The model demonstrates the ability to engage in spontaneous cooperation, leading to increased social welfare beyond mere self-interest, without the necessity for explicit communication. However, this phenomenon also results in low performance in Betraying Games. The model shows limitations in sequential games characterized by more complicated rules. The aggregate score of the model on $\gamma$-Bench is $44.9$.

### v5

# Competing Large Language Models in Multi-Agent Gaming Environments

@@ Abstract
Decision-making is a complex process requiring diverse abilities, making it an excellent framework for evaluating Large Language Models (LLMs). Researchers have examined LLMs’ decision-making through the lens of Game Theory. However, existing evaluation mainly focus on two-player scenarios where an LLM competes against another. Additionally, previous benchmarks suffer from test set leakage due to their static design. We introduce GAMA($\gamma$)-Bench, a new framework for evaluating LLMs’ Gaming Ability in Multi-Agent environments. It includes eight classical game theory scenarios and a dynamic scoring scheme specially designed to quantitatively assess LLMs’ performance. $\gamma$-Bench allows flexible game settings and adapts the scoring system to different game parameters, enabling comprehensive evaluation of robustness, generalizability, and strategies for improvement. Our results indicate that GPT-3.5 demonstrates strong robustness but limited generalizability, which can be enhanced using methods like Chain-of-Thought. We also evaluate 13 LLMs from 6 model families, including GPT-3.5, GPT-4, Gemini, LLaMA-3.1, Mixtral, and Qwen-2. Gemini-1.5-Pro outperforms others, scoring of $69.8$ out of $100$, followed by LLaMA-3.1-70B ($65.9$) and Mixtral-8x22B ($62.4$). Our code and experimental results are publicly available at https://github.com/CUHK-ARISE/GAMABench.
@@ 1 Introduction
We have recently witnessed the advancements in artificial intelligence made by Large Language Models (LLMs), which have marked a significant breakthrough in the field. ChatGPT [^1] https://chat.openai.com/, a leading LLM, has demonstrated its proficiency in a variety of natural language processing tasks, including machine translation Jiao et al. (2023), sentence revision Wu et al. (2023), information retrieval Zhu et al. (2023), and program repair Surameery & Shakor (2023). Beyond the academic sphere, LLMs have entered diverse aspects of our everyday life, such as education Baidoo-Anu & Ansah (2023), legal service Guha et al. (2023), product design Lanzi & Loiacono (2023), and healthcare Johnson et al. (2023). Given their extensive capabilities, evaluating LLMs demands more than simple, isolated tasks. A comprehensive and multifaceted approach is highly in demand to assess the efficacy of these advanced models.
With the broad knowledge encoded in LLMs, their intelligence Liang et al. (2024), and capabilities in general-purpose task solving Qin et al. (2023), a question emerges: Can LLMs assist in everyday decision-making? Many real-world decision-making scenarios can be modeled using Game Theory Koller & Pfeffer (1997). Furthermore, individuals’ ability to achieve Nash equilibrium Nash (1950) reflects their capacity in decision-making Risse (2000). Therefore, many studies have drawn on the principles of game theory Duan et al. (2024); Xie et al. (2024); Xu et al. (2023a), which has several advantages: (1) Scope: Game theory allows for the abstraction of diverse real-life scenarios into simple mathematical models, facilitating a broad range of evaluations. (2) Quantifiability: By examining the Nash equilibrium within these models, we gain a measurable metric for comparing LLMs’ decision-making performance. (3) Variability: The adjustable parameters of these models enable the creation of variant scenarios, enhancing the diversity and robustness of our assessments. However, existing research is often limited to two-player or two-action settings, such as the classical Prisoner’s Dilemma and Ultimatum Game Guo (2023); Phelps & Russell (2023); Akata et al. (2023); Aher et al. (2023); Brookins & DeBacker (2024). Moreover, prior work relies on fixed, classical game settings, increasing the likelihood that LLMs have encountered these scenarios during training, facing the risk of test set leakage. In this paper, we assess LLMs in more complex scenarios involving multiple players, actions, and rounds, across classical game theory scenarios with dynamically adjustable game parameters.
Table 1: Performance (scores) of different LLMs on $\gamma$-Bench.
@@
0613 1106 0125 t-0125 o-0806 1.0 1.5
Guess 2/3 of the Average $41.4_{\pm 0.5}$ $68.5_{\pm 0.5}$ $63.4_{\pm 3.4}$ $91.6_{\pm 0.6}$ $94.3_{\pm 0.6}$ $77.3_{\pm 6.2}$ $95.4_{\pm 0.5}$
El Farol Bar $74.8_{\pm 4.5}$ $64.3_{\pm 3.1}$ $68.7_{\pm 2.7}$ $23.0_{\pm 8.0}$ $70.0_{\pm 22.1}$ $33.5_{\pm 10.3}$ $37.2_{\pm 4.2}$
Divide the Dollar $42.4_{\pm 7.7}$ $70.3_{\pm 3.3}$ $68.6_{\pm 2.4}$ $98.1_{\pm 1.9}$ $95.2_{\pm 0.7}$ $77.6_{\pm 3.6}$ $93.8_{\pm 0.3}$
Public Goods Game $17.7_{\pm 1.7}$ $43.5_{\pm 12.6}$ $38.9_{\pm 8.1}$ $89.2_{\pm 1.8}$ $90.9_{\pm 3.0}$ $68.5_{\pm 7.6}$ $100.0_{\pm 0.0}$
Diner’s Dilemma $67.0_{\pm 4.9}$ $1.4_{\pm 1.3}$ $2.8_{\pm 2.8}$ $0.9_{\pm 0.7}$ $10.7_{\pm 8.3}$ $3.1_{\pm 1.5}$ $35.9_{\pm 5.3}$
Sealed-Bid Auction $10.3_{\pm 0.2}$ $7.6_{\pm 1.8}$ $13.0_{\pm 1.5}$ $24.2_{\pm 1.1}$ $20.8_{\pm 3.2}$ $31.6_{\pm 12.2}$ $26.9_{\pm 9.4}$
Battle Royale $19.5_{\pm 7.7}$ $35.7_{\pm 6.8}$ $28.6_{\pm 11.0}$ $86.8_{\pm 9.7}$ $67.3_{\pm 14.8}$ $16.5_{\pm 6.9}$ $81.3_{\pm 7.7}$
Pirate Game $68.4_{\pm 19.9}$ $69.5_{\pm 14.6}$ $71.6_{\pm 7.7}$ $85.4_{\pm 8.7}$ $84.4_{\pm 6.7}$ $57.4_{\pm 14.3}$ $87.9_{\pm 5.6}$
Overall $42.7_{\pm 2.0}$ $45.1_{\pm 1.6}$ $44.4_{\pm 2.1}$ $62.4_{\pm 2.7}$ $66.7_{\pm 4.7}$ $45.7_{\pm 3.4}$ $69.8_{\pm 1.6}$
(a) Closed-source LLMs: Gemini-1.5-Pro outperforms. For GPT-4: t denotes Turbo and o denotes Omni.
@@
Public Goods Game $19.6_{\pm 1.0}$ $90.6_{\pm 3.6}$ $97.0_{\pm 0.8}$ $27.6_{\pm 11.7}$ $83.7_{\pm 3.5}$ $81.3_{\pm 5.9}$
@@
Sealed-Bid Auction $37.1_{\pm 3.1}$ $15.7_{\pm 2.7}$ $14.7_{\pm 3.2}$ $3.1_{\pm 1.6}$ $13.2_{\pm 3.7}$ $2.5_{\pm 0.7}$
Battle Royale $35.9_{\pm 12.1}$ $77.7_{\pm 26.0}$ $92.7_{\pm 10.1}$ $12.6_{\pm 9.4}$ $36.0_{\pm 21.0}$ $81.7_{\pm 9.6}$
Pirate Game $78.3_{\pm 10.0}$ $64.0_{\pm 15.5}$ $65.6_{\pm 22.3}$ $67.3_{\pm 7.6}$ $84.3_{\pm 8.8}$ $86.1_{\pm 6.4}$
Overall $56.0_{\pm 3.1}$ $65.9_{\pm 3.3}$ $61.8_{\pm 4.7}$ $43.4_{\pm 2.2}$ $62.4_{\pm 2.2}$ $56.7_{\pm 3.4}$
(b) Open-source LLMs: LLaMA-3.1-70B outperforms.
@@
We include eights games and divide them into three categories based on their characteristics. The first category in our framework evaluates LLMs’ ability to make optimal decisions by understanding game rules and recognizing patterns in other players’ behavior. A distinctive characteristic of these games is that individual players cannot achieve higher gains without cooperation, provided that other participants cooperate. Essentially, these games’ Nash equilibrium aligns with maximizing overall social welfare. We name such games as I. Cooperative Games, including (1) Guess 2/3 of the Average, (2) El Farol Bar, and (3) Divide the Dollar. The second category assesses the propensity of LLMs to prioritize self-interest, potentially betraying others for greater gains. In contrast to the first category, games in this category incentivize higher rewards for participants who betray their cooperative counterparts. Typically, the Nash equilibrium in these games leads to reduced social welfare. This category is termed II. Betraying Games, including (4) Public Goods Game, (5) Diner’s Dilemma, (6) Sealed-Bid Auction. Last but not least, we focus specifically on two games characterized by sequential decision-making processes, distinguishing them from the previous six games based on simultaneous decision-making. III. Sequential Games are the (7) Battle Royale and (8) Pirate Game.
Decision-making is a complex task requiring various abilities. Several common ones are evaluated across all games: (1) Perception: the ability to understand situations, environments, and rules, and extends to long-text understanding for LLMs. (2) Arithmetic Reasoning: the ability to quantify real-world options and perform calculations. (3) ToM Reasoning: the Theory of Mind Kosinski (2023); Bubeck et al. (2023); Huang et al. (2024a) refers to the ability to infer others’ intentions and beliefs. (4) Strategic Reasoning: the ability to integrate all available information to arrive at the best decision. Certain games involve specialized abilities, such as K-level reasoning in the “Guess 2/3 of the Average” game and mixed strategy adoption in the “El Farol Bar” game.
@@
In this paper, we instruct ten agents, based on the GPT-3.5 (0125) model, to engage in the eight games, followed by an analysis of the results obtained. Subsequently, we assess the model’s robustness against multiple runs, temperature parameter alterations, and prompt template variations. Further exploration is conducted to ascertain if instructional prompts, such as Chain-of-Thought (CoT) Kojima et al. (2022), enhance the model’s decision-making capabilities. Additionally, the model’s capacity to generalize across diverse game settings is examined. Finally, we evaluate the performance of thirteen LLMs, including GPT-3.5-Turbo (0613, 1106, 0125) OpenAI (2022), GPT-4 (Turbo-0125, 4o-0806) OpenAI (2023), Gemini-1.0-Pro Pichai & Hassabis (2023), Gemini-1.5-Pro Pichai & Hassabis (2024), LLaMA-3.1 (8B, 70B, 405B) Dubey et al. (2024), Mixtral (8x7B, 8x22B) Jiang et al. (2024), and Qwen-2-72B Yang et al. (2024). We compare the performance of different LLMs by creating multiple agents from the same model to participate in the games, then calculate the average performance of these agents. Our contributions include:
  • We provide a comprehensive review and comparison of existing literature on evaluating LLMs using game theory scenarios, as summarized in Table 3. The review includes key aspects such as models, games, temperature settings, and other game parameters, highlighting our emphasis on the multi-player setting and the generalizability of LLMs.
  • Starting from the multi-player setting, we collect eight classical game theory scenarios to measure LLMs’ Gaming Ability in Multi-Agent environments, and implement our framework, GAMA($\gamma$)-Bench. It enables dynamic game scene generation with diverse profiles, offering unlimited scenarios to assess LLM generalizability while minimizing test set leakage risk.
@@
  • We apply $\gamma$-Bench to thirteen LLMs to provide an in-depth analysis of their performance in multi-agent gaming scenarios, indicating their potential as assistants in decision-making process.
2 Introduction to Games
@@
We collect eight games well studied in Game Theory and propose $\gamma$-Bench, a framework with multi-player, multi-round, and multi-action settings. Notably, $\gamma$-Bench allows the simultaneous participation of both LLMs and humans, enabling us to evaluate LLMs’ performance when playing against humans or fixed strategies. This section details each game with their classical settings (parameters).
@@ 2.1 Cooperative Games
Initially introduced by Ledoux (1981), the game involves players independently selecting an integer between 0 and 100 (inclusive). The winner is the player(s) choosing the number closest to two-thirds of the group’s average. A typical initial strategy might lead players to assume an average of 50, suggesting a winning number around $50\times\frac{2}{3}\approx 33$. However, if all participants adopt this reasoning, the average shifts to 33, thereby altering the winning number to approximately 22. The game has a Pure Strategy Nash Equilibrium (PSNE) where all players selecting zero results in a collective win.
@@
Proposed by Arthur (1994) and Huberman (1988), this game requires players to decide to either visit a bar for entertainment or stay home without communication. The bar, however, has a limited capacity and can only accommodate part of the population. In a classical scenario, the bar becomes overcrowded and less enjoyable if more than 60% of the population decides to go there. Conversely, if 60% or fewer people are present, the experience is more enjoyable than staying home. Imagine that if everyone adopts the same pure strategy, i.e., either everyone going to the bar or everyone staying home, then the social welfare is not maximized. Notably, the game lacks a PSNE but presents an Mixed Strategy Nash Equilibrium (MSNE), where the optimal strategy involves going to the bar with a 60% probability and staying home with a 40% probability.
@@ 2.3 Sequential Games
3 GAMA-Bench Scoring Scheme
@@
Key Findings: The model’s decisions are mainly influenced by the outcomes of the preceding round rather than deriving from the reasoning of the optimal strategy. Although initially demonstrating suboptimal performance, the model can learn from historical data and enhance its performance over time. A larger fluctuation is observed in games that are difficult to optimize from historical data, such as the El Farol Bar game. The model demonstrates the ability to engage in spontaneous cooperation, leading to increased social welfare beyond mere self-interest, without the necessity for explicit communication. However, this phenomenon also results in low performance in Betraying Games. The model shows limitations in sequential games with more complicated rules. The aggregate score of the model on $\gamma$-Bench is $45.9$.